OpenStudy (anonymous):

An unbiased coin is tossed five times. Find the probability of the given event. (Round your answer to four decimal places.) The coin lands heads more than once.

OpenStudy (anonymous):

Chance that the 1st ball is blue = 5/8 The 2nd = 4/7 The 3rd = 3/6 the 4rth = 2/5 multiply these numbers 1/14 that all balls are blue Decimal = 0.071

OpenStudy (anonymous):

5/5....1???